# Towards Fine-Grained Citation Evaluation in Generated Text: A Comparative Analysis of Faithfulness Metrics

### v1

@@ 4.3 LLM-Based Metrics
5 Experiments
5.1 Datasets
Data Statistics
In the experiments, we employ the dataset Liu et al. (2023) collected from generative search engines (e.g., BingChat) as our evaluation benchmark, referred to as GenSearch. This dataset comprises a total of $12,681$ statement-citation pairs. For each pair, human assessors categorize the citation into one of three categories of support levels: full, partial, or no support. The details of data statistics are shown in Table 1. Notably, for citations classified under the full or partial support categories, human assessors additionally extract explicit evidence from the citation that supports the associated statement.
Data Processing
While the GenSearch dataset aligns well with our research objectives, we encounter a significant challenge: the extensive length of most citations within the dataset. These citations often comprise a web document with thousands of words, far exceeding the maximum input capacity of most faithfulness metrics, which is limited to $512$ words. This limitation necessitates input truncation, potentially compromising the reliability of faithfulness metrics. To mitigate this issue, we adopt a strategy similar to previous studies Zha et al. (2023). Specifically, we segment each cited document into shorter text chunks, with a maximum length of $150$ words per chunk. These text chunks, along with their corresponding statements, serve as the inputs for faithfulness metrics to predicted metric scores. Furthermore, to determine human judgments for the text chunks, we employ the Jaccard similarity index to identify text chunks containing human-annotated evidence, classifying them as either fully or partially supporting text chunks.
5.2 Evaluation Metrics
@@ 7 Discussions
Overall, our results indicate that the evaluation protocols are complementary and should be integrated for a comprehensive assessment of metrics. Based on the evaluation results, we further propose the following practical recommendations to develop more effective metrics for automated citation evaluation: 1) Development of training resources:motivated by the observation that the best-performing metrics still struggle with identifying partial support, we recommend the development of training resources that include fine-grained support level annotations. These resources could significantly enhance the metrics’ fine-grained sensitivity to varying support levels; 2) Introduction of contrastive learning:to improve the robustness of metrics in post-hoc retrieval scenarios, we recommend fine-tuning metrics using contrastive learning frameworks. This method has demonstrated effectiveness across various information retrieval tasks Izacard et al. (2022). 3) Development of more explainable metrics:traditional faithfulness metrics often only provide final scores without sufficient explainability. This limitation hinders a deeper understanding of the models’ behavior. Therefore, it is crucial to develop more explainable faithfulness metrics, potentially using large language models (LLMs).
@@ 8 Conclusion
In this work, we consider a citation that explicitly contains human-annotated evidence as the fully supporting citation for each statement. However, for some complex statements, their evidence is distributed among multiple citations. For instance, about $2\%$ statements on the GenSearch dataset require multiple citations to be fully supported. Also, we focus on statement-level citation evaluation. Since answer-level citation evaluation is much more complicated and requires proper aggregation methods, we leave this exploration as future work. We do not evaluate QA-based faithfulness metrics as a recent study shows that such metrics have some fundamental issues, such as failing to localize errors Kamoi et al. (2023). However different findings could be explored with QA-based metrics.
@@ Appendix A Details of Prompts
Details of Prompts are shown in Table 5.
@@
Table 5: Prompts for discrete and continuous scoring methods.

### v2

@@ 7 Discussions
Overall, our results across three evaluation protocols indicate that the evaluation protocols are complementary and should be integrated for a comprehensive assessment of metrics. Based on the evaluation results, we further propose the following practical recommendations to develop more effective metrics for automated citation evaluation: 1) Development of training resources:motivated by the observation that the best-performing metrics still struggle with identifying partial support, we recommend the development of training resources that include fine-grained support level annotations. These resources could significantly enhance the metrics’ fine-grained sensitivity to varying support levels; 2) Introduction of contrastive learning:to improve the robustness of metrics in post-hoc retrieval scenarios, we recommend fine-tuning metrics using contrastive learning frameworks. This method has demonstrated effectiveness across various information retrieval tasks Izacard et al. (2022). 3) Development of more explainable metrics:traditional faithfulness metrics often only provide final scores without sufficient explainability Xu et al. (2023). This limitation hinders a deeper understanding of the models’ behavior. Therefore, it is crucial to develop more explainable faithfulness metrics, potentially using large language models (LLMs).
@@ 8 Conclusion
In this work, we consider a citation that explicitly contains human-annotated evidence as the fully supporting citation for each statement. However, for some complex statements, their evidence is distributed among multiple citations. For instance, about $2\%$ statements on the GenSearch dataset require multiple citations to be fully supported. Also, we focus on statement-level citation evaluation. Since answer-level citation evaluation is much more complicated and requires proper aggregation methods, we leave this exploration as future work. We do not evaluate QA-based faithfulness metrics as a recent study shows that such metrics have some fundamental issues, such as failing to localize errors Kamoi et al. (2023). However different findings could be explored with QA-based metrics. Finally, there is another line of work, in which they break down a complex statement into multiple atomic facts.
@@
Acknowledgments
This research was supported by the China Scholarship Council (CSC) under grant number 202008440470. The views expressed in the content belong solely to the authors and may not reflect the perspectives or endorsements of their respective employers or sponsors.
@@ Appendix A Details of Prompts
Details of GPT-3.5 prompts used in the experiments are shown in Table 5.
@@
Table 5: Detailed prompts for discrete and continuous scoring methods.